OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6

OpenStudy (anonymous):

First you have to minus 3 from both sides

OpenStudy (anonymous):

First distribute -5*x=-5x -5*2=-10 On the other side 3*x=3x 3*1= 3 rewrite the equation it will be 3-5x-10= 3x+3+6 try to bring all the variabkes on one side take 5x and add it to 3x which makes it 8x that makes 5x cancel out now rewrite the equation it would be 3-10 =8x+9 -7=8x+9 subtract -9 from each side which leaves you with -16=8x so the answer would be -2=x